A manufacturer of bird feeder cut spaces from tube that 5 1/4 inches long. How many spaces can be cut from tube if each spacer must be 3/4 inch thick.
Answer by nerdybill(7384)

A manufacturer of bird feeder cut spaces from tube that 5 1/4 inches long. How many spaces can be cut from tube if each spacer must be 3/4 inch thick.
.
5 1/4 = 5 + 1/4 = 20/4 + 1/4 = 21/4
.
(21/4)/(3/4)
(21/4)*(4/3)
(21/1)*(1/3)
(21/3)
7